A motor boat is heading due north and crosses a wide river with a speed of 4m/s relative to the water. the water in the river has a uniform speed of 3m/s due east relative to the earth . determine the direction of the motion of the motor boat relative to the earth
1
Expert's answer
2020-04-10T08:51:28-0400

Notations

  • Velocity of the boat relative to the river: V(b.r)
  • Velocity of the river relative to earth v(r,e)
  • Velocity of the boat relative to earth: V(b,e)
  • Angle between V(b,e) & V(r,e) velocity vectors "\theta"


Calculations



  • Considering the vector triangle which could be formed by all the velocity vectors the boat is exposed to (Refer the sketch attached),

"\qquad
\begin{aligned}
\small 
\end{aligned}" "\qquad
\begin{aligned}
\small \tan{\theta}&=\small \frac{V_{b,r}}{V_{r,e}}\\
\small \theta&= \small \tan^{-1}(\frac{4\,ms^{-1}}{3\,{ms^{-1}}})\\
\small &= \small \bold{53.13^{\degree}} 
\end{aligned}"

Therefor, the direction of the boat is "\small \bold {53.13^{\degree}}" to north from east.


  • Additionally its velocity,

"\qquad \qquad
\begin{aligned}
\small V^2_{b,e}& = \small V^2_{b,r}\,+\,V^2_{r,e}\\
\small V_{b,e}&=\small \sqrt{(4ms^{-1})^2\,+\,(3ms^{-1})^2}\\
\small&=\small \bold{5ms^{-1}}
\end{aligned}"